OpenStudy (anonymous):

-n + 7 - 2n + 5n -3 = -6 Can someone show me the long term

OpenStudy (anonymous):

hmmmm. could be....

OpenStudy (anonymous):

long term? if u wanna find n, then, -n+7-2n+5n-3=-6 that implies +4n-2n+4=-6 that implies, 2n=-10, and n=-10/2 = -5

OpenStudy (mimi_x3):

-n - 2n + 5n +7-3 +6=0 2n +10=0 2n = -10 n = -5

OpenStudy (anonymous):

Thanks

OpenStudy (mimi_x3):

yw(:

OpenStudy (anonymous):

welcum..